OpenStudy (anonymous):

A radioactive gas with a half life of 2.4 days accidentally leaks into a room. If the radiation level is now 25% above a safe amount for how long should the room be emptied?

OpenStudy (anonymous):

1.00 = (1.25)e^(-0.693)(2.4)t 1.00/1.25 = e^(-1.6632)t 0.8 = e^(-1.6632)t t = 0.135 days = 3 hrs 15 min

OpenStudy (anonymous):

@amistre64 please confirm

OpenStudy (amistre64):

im not sure what a safe amount of radiation is to be able to determine that

OpenStudy (anonymous):

the problem says that's all the information i need to solve. that's how i got confused.

OpenStudy (amistre64):